OpenStudy (anonymous):

\sin \theta = \frac{-\sqrt{3}}{2}. Find all angles between 0 and 2\pi satisfying the condition If someone could explain how to do this problem I'd be very grateful

OpenStudy (anonymous):

\[\sin \theta = \frac{-\sqrt{3}}{2}.\]